### v3

@@ 3 Exploiting the antistalling rule for general bounds
If these exists a constant $\lambda>0$ such that
@@
$(n-m)\Big{\lceil}\lambda\log_{e}\frac{m\Delta}{\delta}\Big{\rceil}\,.$
Given (1), apply the simplex algorithm with the antistalling pivot rule described in the previous section. In particular, at a general iteration $t$ of the algorithm, let $B^{t},x^{t}$, and $y^{t}$ be respectively the basis, the basic solution, and the improving feasible direction considered by the algorithm. Let $N^{t}:=[n]\setminus B^{t}$. If $x^{t}\neq x^{t-1}$ (i.e., we encounter $x^{t}$ for the first time, as a result of a non-degenerate pivot) let $y^{t}:=x^{\star}-x^{t}$. Observe that, once this is specified, the improving feasible direction is determined by the antistalling rule for all remaining degenerate pivots at $x^{t}$.
@@
$\begin{array}[]{rcl}c^{T}x^{t+k}-c^{T}x^{\star}&=&c^{T}x^{t}-c^{T}x^{\star}\\ &=&-c^{T}y^{t}\\ &\leq&-c^{T}y^{t+k}\\ &=&-\overline{c}^{T}_{N^{t+k}}y^{t+k}_{N^{t+k}}\\ &=&\sum_{i\in N^{t+k}}-\overline{c}^{T}_{N^{t+k},i}y^{t+k}_{i}\\ &=&\sum_{i\in N^{t+k}\mid y^{t+k}_{i}>0}-\overline{c}^{T}_{N^{t+k},i}y^{t+k}_{i}\\ &\leq&(n-m)\Delta^{t+k}_{\overline{c}}\Delta\,,\end{array}$ (6)
where we used Lemma 2(a) for the first inequality, and (2) for the third equality. The last equality follows from $y^{t+k}_{i}\geq 0,i\in N^{t+k}$ which is implied by feasibility of $y^{t+k}$ at $x^{t+k}$. The last inequality in turn follows from $\max_{\{i\in N^{t+k}\mid y^{t+k}_{i}>0\}}-\overline{c}_{N^{t+k},i}=\Delta^{t+k}_{\overline{c}}$, the fact that $|N^{t+k}|=n-m$, and since $y^{t+k}_{i}=y^{t}_{i}=x^{\star}_{i}-x^{t}_{i}\leq\Delta$ holds for any $i\in N^{t+k}\cap\,\mathrm{supp}(y^{t+k})$, using Lemma 2(b).
@@
Now, combining Lemma 1 and Lemma 3 allows to conclude that the simplex algorithm with the antistalling pivot rule described in Section 2 encounters at most $(n-m)\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$ distinct basic feasible solutions. Since Theorem 1.1 and Remark 1 in turn bound the number of consecutive non-degenerate steps by $\min\{n-m-1,m-1\}$ (again, due to the choice of the improving feasible direction), combining it with the latter result yields a bound on the total number of pivots required to reach an optimal vertex. However, this bound does not take into account the number of (degenerate) pivots that might have to be performed at an optimal vertex to encounter a basis satisfying the optimality criterion. This is due to the fact that the antistalling pivot rule requires an improving feasible direction, which we do not have at an optimal vertex. Hence, we have to handle this case separately.
Let $B$ and $B^{\star}$ be a non-optimal and an optimal basis, both associated with an optimal basic feasible solution $x^{\star}$ of (1). Assume $B$ is the basis of $x^{\star}$ obtained by the antistalling pivot rule. Since $B$ is not optimal, there exists $f\in N$ with $\overline{c}_{N,f}=c^{T}z<0$ with $z$ as in (3). Observe that there exists $i\in B\cap N^{\star}$ with $\overline{A}_{if}>0$ and $x_{i}=0$: otherwise, all coordinates of $z_{N^{\star}}$ would be non-negative and hence $z\in C_{N^{\star}}:=\{x\in\mathrm{\hskip 0.28453ptIR}^{n}\mid Ax=0,x_{N^{\star}}\geq 0\}$. The latter however contradict the fact that all extreme rays of the above cone $C_{N^{\star}}$ have non-negative scalar products $\overline{c}_{N^{\star}}$ with $c$ due to optimality of $B^{\star}$. Hence one could perform a simplex pivot on $B$ with entering variable $f$ and leaving variable $i$. Let $B^{\prime}:=\big{(}B\setminus\{i\}\big{)}\cup\{f\}$ and $N^{\prime}:=\big{(}N\setminus\{f\}\big{)}\cup\{i\}$. Note that $\overline{c}_{N^{\prime},i}\geq 0$ and $i\in N^{\prime}\cap N^{\star}$. Either $B^{\prime}$ is a optimal basis and we stop, or there exists $j\in N^{\prime}\setminus\{i\}$ with $\overline{c}_{N^{\prime},j}<0$. In the latter case, however, we can enforce the constraint $x_{i}=0$ by removing the variable $x_{i}$ together with the corresponding column of $A$ and entry $c_{i}$ of $c$ from (1). By doing so we obtain a new LP with the number of variables smaller by one that has $B^{\prime}$ and $B^{\star}$ as a non-optimal and an optimal basis, respectively, since $\overline{c}_{N^{\prime}\setminus\{i\},j}<0$ and $\overline{c}_{N^{\star}\setminus\{i\}}\geq 0$. We can set $B=B^{\prime}$ and repeat the process. Since at each iteration a variable with index from $N^{\star}$ gets removed, after at most $n-m$ iterations $B=B^{\star}$. Since the number of degenerate pivots at an optimal vertex is bounded by $n-m$, we can state the following result.
@@
$\min\{n-m,m\}(n-m)\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$
@@
In general, the value of $\delta$ is NP-hard to compute [24]. However, observe that for integral $A$ and $b$ (which can be assumed without loss of generality for rational LPs), one can bound $\Delta\leq||b||_{1}\Delta_{A}$ and $\delta\geq\frac{1}{\Delta_{A}}$ due to Cramer’s rule, where $\Delta_{A}$ is the largest absolute value of the determinant of a square submatrix of $A$. Then the next statement a straightforward corollary of the above theorem.
@@ 3.1 Application to combinatorial LPs
One can compute the corresponding pivot sequence for the problems mentioned in (a)-(d) by running the simplex with the antistalling pivot rule. However, one needs to presolve the LP and obtain an optimal basic solution in order to use it to guide the antistalling pivot rule.
@@ 4 Computational experiments
In this section we report some computational experiments that we conducted to evaluate the performance of the antistalling pivot rule when applied to actual linear programs.
For the implementation, we used the python package CyLP [41] which wraps COIN-OR’s CLP solver [9] and provides tools for implementing a preferred pivot rule in python to substitute CLP’s built-in ones. For the tests, we used the benchmark Netlib LP dataset [30] containing 93 LPs of diverse dimensions and sparsity. All our experiments were conducted on a laptop with Intel Core i7-13700H 2.40GHz CPU and 32GB RAM.
For each of the aforementioned LPs, we ran the simplex algorithm with our antistalling pivot rule, and in addition with other 5 well-known pivot rules (which [41] provides implementations for). These are:
@@
  • Steepest edge: the entering variable is the one which yields a direction $z$ maximizing $\frac{-c^{T}z}{||z||_{1}}$ (i.e., maximizing the improvement normalized according to the 1-norm);
@@
We tracked the number of pivots required by each of the above rules. For our antistalling pivot rule, we guided it using the feasible direction $y=x^{\star}-x$ at any non-optimal vertex $x$ of the LP, for a pre-computed optimal basic solution $x^{*}$ (in particular, the one founded by Dantzig’s pivot rule). Intuitively, this represents the best direction that can possibly guide our rule, as it is the direction leading immediately to an optimal solution. Our objective with the computational experiments is to see whether this choice translates into few pivots in practice, since $y$ might change during the degenerate steps and therefore in reality we do not have control on the actual edge-direction we end up moving along.
Among the 93 instances in the dataset, there were 84 LPs for which each of the pivot rules was able to find an optimal solution with 30 minutes of timeout. We restrict our report to these LPs. For the sake of graphical representation, we further divided these 84 problems into 2 groups of 42 LPs each, by considering the maximum number of iterations required by any of the 6 tested pivot rules. In particular, the first group contains 42 LPs for which all pivot rules were able to arrive to an optimal solution within 2369 pivots. The second group contains the remaining 42 LPs. Figure 2 shows the results.
Figure 2: Both the upper and the lower plots show the number of iterations required to solve the LPs by each of the 6 considered pivot rules. The 42 LPs in the upper plot are the ones solved within 2369 pivots by all pivot rules. The remaining 42 LPs are considered in the lower plot. Each of the so-called violin plot visualises the distribution of the corresponding 42 numbers. White dots depict means, thick black dashes represent the intervals between 0.25- and 0.75-quantiles.
@@
The computational experiments show that the antistalling pivot rule, guided by a known optimal solution, actually performs quite well, and most often manages to find a relatively short sequence of pivots to an optimal basis compared to other pivot rules. We highlight in particular the problem 25fv47, where any other pivot rule required at least 7892 pivots whereas the antistalling only needed 1468. On the other hand, there are instances where the antistalling pivot rule actually showed the worse performance compared to all its competitors. Those are the LPs fit2p and seba: 69711 and 976 pivots against at most 61925 and 873, respectively.
@@ 5 Final Remarks
Next, we would like to connect the above question to the recent work [20]. Given a basis $B$ defining a degenerate vertex $v$, one could imagine to separate $v$ from all its neighbouring vertices by a hyperplane: this way, one obtain a pyramid (with $v$ at its apex). Suppose this pyramid is full-dimensional and is described by some $Ax\leq b$: then, the above question is equivalent to perturbing the side facets of the pyramid to allow for a short path from the new apex defined by the given basis $B$, to any of the vertices in the bottom of the pyramid. This can be done relying on the techniques used for the construction of a so-called rock extension [20] (when viewing the base of the pyramid as the $d-1$-dimensional original polytope, and the pyramid as its $d$-dimensional extension). Such rock extension has been proven to have the property that the apex is connected to any bottom vertex by a path of at most linear length. However these paths are not necessarily monotone for a given linear objective function, and moreover, the construction of the rock extension requires not only perturbing the right-hand-side $b$ but the left-hand-side $A$ as well.
@@
Futhermore, as it was mentioned in Section 2, in order for our bound to hold, the only requirement we need to impose when performing a pivot is that the entering variable lies in the support of $y$ (besides of course having negative reduced cost). In order to obtain the results from Section 3, where both degenerate and non-degenerate pivots have to be taken into account, we chose the entering variable to be the one with the most negative reduced cost, i.e., we used Dantzig’s rule just restricted to the coordinates in $\mathrm{supp}(y)$. Of course, one could think of analyzing the performance of the antistalling pivot rule with other choices of entering coordinates from $\mathrm{supp}(y)$, e.g., according to steepest edge or shadow vertex. Variations of these latter two rules, in particular, have been shown to play a key role in the context of 0/1 polytopes [4].

### v4

@@ 3 Exploiting the antistalling rule for general bounds
If there exists a constant $\lambda>0$ such that
@@
$(n-m)\Big\lceil\lambda\ln\frac{m\Delta}{\delta}\Big\rceil\,.$
Given (1), apply the simplex algorithm with the antistalling pivot rule described in the previous section. In particular, at a general iteration $t$ of the algorithm, let $B^{t},x^{t}$, and $y^{t}$ be the basis, the basic solution, and the improving feasible direction considered by the algorithm, respectively. Let $N^{t}:=[n]\setminus B^{t}$. If $x^{t}\neq x^{t-1}$ (i.e., we encounter $x^{t}$ for the first time, as a result of a non-degenerate pivot) let $y^{t}:=x^{\star}-x^{t}$. Observe that, once this is specified, the improving feasible direction is determined by the antistalling rule for all remaining degenerate pivots at $x^{t}$.
@@
where we used Lemma 2(a) for the first inequality, and (2) for the third equality. The last equality follows from $y^{t+k}_{i}\geq 0,i\in N^{t+k}$ which is implied by feasibility of $y^{t+k}$ at $x^{t+k}$. The last inequality follows from $\max_{\{i\in N^{t+k}\mid y^{t+k}_{i}>0\}}-\overline{c}_{N^{t+k},i}=\Delta^{t+k}_{\overline{c}}$, the fact that $|N^{t+k}|=n-m$, and because $y^{t+k}_{i}=y^{t}_{i}=x^{\star}_{i}-x^{t}_{i}\leq\Delta$ holds for any $i\in N^{t+k}\cap\,\mathrm{supp}(y^{t+k})$, using Lemma 2(b).
Let $x^{t+k}_{f}$ denote the entering variable at $(t+k)\textsuperscript{th}$ iteration. Note that $x^{t+k+1}_{f}\neq 0$ since $x^{t+k}\neq x^{t+k+1}$. Then,
@@
Now, combining Lemma 1 and Lemma 3 allows to conclude that the simplex algorithm with the antistalling pivot rule described in Section 2 encounters at most $(n-m)\lceil\frac{(n-m)\Delta}{\delta}\ln\big(m\frac{\Delta}{\delta}\big)\rceil$ distinct basic feasible solutions. Since Theorem 2.1 bounds the number of consecutive non-degenerate steps by $\min\{n-m-1,m-1\}$ (again, due to the choice of the improving feasible direction), combining it with the latter result yields a bound on the total number of pivots required to reach an optimal vertex. However, this bound does not take into account the number of (degenerate) pivots that might have to be performed at an optimal vertex to encounter a basis satisfying the optimality criterion. This is due to the fact that the antistalling pivot rule requires an improving feasible direction, which we do not have at an optimal vertex. Hence, we have to handle this case separately.
Lemma 4
Let $B$ and $B^{\star}$ be a non-optimal and an optimal basis, respectively, both associated with an optimal basic feasible solution $x^{\star}$ of (1). We can obtain $B^{\star}$ from $B$ by performing at most nmn-m degenerate simplex pivots.
Proof
Since $B$ is not optimal, there exists $f\in N$ with $\overline{c}_{N,f}=c^{T}z<0$ with $z$ as in (3). Observe that there exists $i\in B\cap N^{\star}$ with $\overline{A}_{if}>0$ and $x_{i}=0$: otherwise, all coordinates of $z_{N^{\star}}$ would be non-negative and hence $z\in C_{N^{\star}}:=\{x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}\mid Ax=0,x_{N^{\star}}\geq 0\}$. The latter however contradicts the fact that all extreme rays of the above cone $C_{N^{\star}}$ have non-negative scalar products $\overline{c}_{N^{\star}}$ with $c$ due to optimality of $B^{\star}$. Hence one could perform a simplex pivot on $B$ with entering variable $f$ and leaving variable $i$. Let $B^{\prime}:=\big(B\setminus\{i\}\big)\cup\{f\}$ and $N^{\prime}:=\big(N\setminus\{f\}\big)\cup\{i\}$. Note that $\overline{c}_{N^{\prime},i}\geq 0$ and $i\in N^{\prime}\cap N^{\star}$. Either $B^{\prime}$ is a optimal basis and we stop, or there exists $j\in N^{\prime}\setminus\{i\}$ with $\overline{c}_{N^{\prime},j}<0$. In the latter case, however, we can enforce the constraint $x_{i}=0$ by removing the variable $x_{i}$ together with the corresponding column of $A$ and entry $c_{i}$ of $c$ from (1). By doing so we obtain a new LP with the number of variables smaller by one that has $B^{\prime}$ and $B^{\star}$ as a non-optimal and an optimal basis, respectively, since $\overline{c}_{N^{\prime}\setminus\{i\},j}<0$ and $\overline{c}_{N^{\star}\setminus\{i\}}\geq 0$. We can set $B=B^{\prime}$ and repeat the process. Since at each iteration a variable with index from $N^{\star}$ gets removed, after at most nmn-m iterations $B=B^{\star}$. ∎
Since the number of degenerate pivots at an optimal vertex is bounded by nmn-m by the above lemma, we can state the following result.
@@
$\min\{n-m,m\}(n-m)\lceil\frac{(n-m)\Delta}{\delta}\ln\big(m\frac{\Delta}{\delta}\big)\rceil$
@@
[…]
@@ 3.1 Application to combinatorial LPs
One can compute the corresponding pivot sequence for the problems mentioned in (a)-(d) by running the simplex with the antistalling pivot rule. However, one needs to first solve the LP and obtain an optimal basic solution in order to use it to guide the antistalling pivot rule.
@@ 4 Computational experiments
In this section we report some computational experiments that we conducted to evaluate the performance of the antistalling pivot rule when applied to actual linear programs. The code is now available at https://github.com/kirill-kukharenko/antistalling.
For the implementation, we used the python package CyLP [44] which wraps COIN-OR’s CLP solver [10] and provides tools for implementing a preferred pivot rule in python to substitute CLP’s built-in ones. We used two different testsets of diverse dimensions and sparsity. The first one is the benchmark Netlib LP dataset [32], containing 108 LPs. The second one contains the root LPs of the benchmark set of MIPLIB 2017 [17], and is composed of 240 instances. All our experiments were conducted on a laptop with Intel Core i7-13700H 2.40GHz CPU and 32GB RAM.
For each of the aforementioned LPs, we ran the simplex algorithm with our antistalling pivot rule, and in addition with other 6 well-known pivot rules (which [44] provides implementations for). These are:
@@
[…]
@@
  • Positive edge rule: the entering variable is selected with a heuristic that tries to identify non-degenerate pivots with high probability [34].
@@
We tracked the number of pivots required by each of the above rules to solve the LPs to optimality (we did not perturb the instances). For our antistalling pivot rule, we guided it using the feasible direction $y=x^{\star}-x$ at any non-optimal vertex $x$ of the LP, for a pre-computed optimal basic solution $x^{*}$ (in particular, the one found by Dantzig’s pivot rule). Intuitively, this represents the best direction that can possibly guide our rule, as it is the direction leading immediately to an optimal solution. Our objective with the computational experiments is to see whether this choice translates into few pivots in practice, since $y$ might change during the degenerate steps and therefore in reality we do not have control on the actual edge-direction we end up moving along.
We remark that we are here interested in the number of pivots required by various pivot rules to solve LPs to optimality, and not in the computational time needed to do so. In fact, the CyLP package connects to the solver through a relatively slow Python interface, in contrast to implementing the rules directly into a solver. Therefore, the efficiency of the pivot rule implementations, and hence the number of problems solved by each rule within the time/memory limit, can be heavily affected by this overhead. For a fair comparison, in this section we then only included the instances solved to optimality by all respective pivot rules. However, for completeness, detailed tables with all the results of the computational experiments, for both Netlib and MIPLIB testsets, can be found in Appendices 0.A and 0.B, respectively.
Our first report deals with 84 out of 108 Netlib problems, for which each of the aforementioned 7 pivot rules was able to find an optimal solution within the time limit of 30 minutes and the memory limit of 32GB RAM. For the results, see Figure 2. Similarly, for MIPLIB this restriction leaves us with 117 out of 240 problems. See Figure 3 for the results.
Figure 2: Netlib dataset results. Data are presented using standard box plots, which display the distribution of a dataset based on its quartiles. Specifically, the box extends from the first quartile (Q1) to the third quartile (Q3), representing the middle $50\%$ of the data. The thin vertical lines (whiskers) extend from the box to the most extreme data points lying within 1.5 times the interquartile range (Q3-Q1). Points beyond the whiskers are shown individually as outliers (fliers). Orange horizontal lines denote the medians: 551, 659, 648, 739, 2027, 1890, and 2049, respectively.
Figure 3: MIPLIB dataset results. Orange lines denote the medians: 1376, 1633, 1493, 2335, 7374, 7829, and 8641, respectively.
@@
The computational experiments show that the antistalling pivot rule, guided by a known optimal solution, actually performs quite well, especially on the Netlib instances. Most often it manages to find a relatively short sequence of pivots to an optimal basis compared to the other pivot rules. We highlight for example the problems 25fv47 from Netlib and istanbul-no-cutoff.mps from MIPLIB, where the antistalling pivot rule required less than one fourth of the number of iterations of any other rule. On the other hand, there are instances, e.g., fit2p from Netlib and ic97_potential.mps from MIPLIB, where the antistalling pivot rule actually showed the worst performance compared to all its competitors.
Overall, the results seem to confirm that guidance by an optimal solution indeed often translates into a small number of pivots, motivating further theoretical studies on how such a direction could be approximately followed.
@@ 5 Final Remarks
Next, we would like to connect the above question to the recent work [23]. Given a basis $B$ defining a degenerate vertex $v$, one could imagine to separate $v$ from all its neighbouring vertices by a hyperplane: this way, one would obtain a pyramid (with $v$ at its apex). Suppose this pyramid is full-dimensional and is described by some $Ax\leq b$: then, the above question is equivalent to perturbing the side facets of the pyramid to allow for a short path from the new apex defined by the given basis $B$, to any of the vertices in the bottom of the pyramid. This can be done relying on the techniques used for the construction of a so-called rock extension [23] (when viewing the base of the pyramid as the d1d-1-dimensional original polytope, and the pyramid as its $d$-dimensional extension). Such rock extension has been proven to have the property that the apex is connected to any bottom vertex by a path of at most linear length. However these paths are not necessarily monotone for a given linear objective function, and moreover, the construction of the rock extension requires not only perturbing the right-hand-side $b$ but the left-hand-side $A$ as well.
@@
Furthermore, as it was mentioned in Section 2, in order for our bound to hold, the only requirement we need to impose when performing a pivot is that the entering variable lies in the support of $y$ (besides of course having negative reduced cost). In order to obtain the results from Section 3, where both degenerate and non-degenerate pivots have to be taken into account, we chose the entering variable to be the one with the most negative reduced cost, i.e., we used Dantzig’s rule just restricted to the coordinates in $\mathrm{supp}(y)$. Of course, one could think of analyzing the performance of the antistalling pivot rule with other choices of entering coordinates from $\mathrm{supp}(y)$, e.g., according to steepest edge or shadow vertex. Variations of these latter two rules, in particular, have been shown to play a key role in the context of 0/1 polytopes [4].